OpenStudy (anonymous):

If a line passes through the points (1,3) and (-2,-9), the equation of the line is y - 3 = 4(x - ___).

OpenStudy (hysenberg):

equation of line passing through (x1,y1) and (x2,y2) is \[\frac{ x-x1 }{ y-y1 }=\frac{ x1-x2 }{ y1-y2 }\]